find the range of aprojectile launched at an angle of 45ยฐ when an intial velocity of 25m/s.

1
Expert's answer
2021-12-07T09:51:41-0500

"l=\frac{v_0^2\sin2\alpha}{2g}=\frac{25^2\cdot\sin(2\cdot45°)}{2\cdot9.8}\approx32\ (m)" . Answer